# Length-Induced Embedding Collapse in Transformer-based Models

### v1

@@ Abstract
Text embeddings enable various applications, but their performance deteriorates on longer texts. In this paper, we find that the performance degradation is due to a phenomenon called Length Collapse, where longer text embeddings collapse into a narrow space. This collapse results in a distributional inconsistency between embeddings of different text lengths, ultimately hurting the performance of downstream tasks. Theoretically, by considering the self-attention mechanism inherently functions as a low-pass filter, we prove that long sequences increase the attenuation rate of the low-pass filter effect of the self-attention mechanism. With layers going deeper, excessive low-pass filtering causes the token signals to retain only their Direct-Current (DC) component, which means the input token feature maps will collapse into a narrow space, especially in long texts. Based on the above analysis, we propose to mitigate the undesirable length collapse limitation by introducing a temperature in $\mathrm{softmax}(\cdot)$, which achieves a higher low-filter attenuation rate. The tuning-free method, called TempScale, can be plugged into multiple transformer-based embedding models. Empirically, we demonstrate that TempScale can improve existing embedding models especially on long text inputs, bringing up to 0.53% performance gains on 40 datasets from Massive Text Embedding Benchmark (MTEB) and 0.82% performance gains on 4 datasets from LongEmbed, which specifically focuses on long context retrieval.
@@ 1 Introduction
Text embeddings—dense vectors that preserve the semantic information of given texts—have become fundamental to many downstream natural language processing (NLP) applications, including text analysis (Aggarwal & Zhai, 2012; Angelov, 2020), question answering (Tan et al., 2023; Xu et al., 2024), web search (Zhao et al., 2023; Yates et al., 2021), and retrieval-augmented generation (Gao et al., 2023; Fan et al., 2024). Typically, embeddings are generated by pre-trained language models (PLMs), which produce fixed-dimensional embeddings regardless of the input text length. In practice, we expect PLMs to perform consistently on texts of varying lengths in any downstream applications.
Unfortunately, we observe that popular transformer-based embedding models perform poorly on longer texts. As shown in Figure 1(a), using the classification task on the IMDB dataset from the Massive Text Embedding Benchmark (MTEB) (Muennighoff et al., 2023) leaderboard as an example, we evaluate the performance of mainstream embedding models on test sets grouped by different text lengths. The experimental results reveal that models of different capabilities and context window sizes consistently exhibit performance degradation as text length increases. For instance, the BGE (Xiao et al., 2023) model’s classification accuracy drops significantly from 75.6% in the length range [0, 100) tokens to 59.0% in the range [400, 500) tokens, indicating a substantial decline of 16.6% points.
@@
We attribute this performance degradation to a biased behavior of embedding models: embeddings of longer texts tend to cluster together, a phenomenon we term as length collapse. To verify this, we conduct controlled experiments depicted in Figures 1(b) and 1(c). Figure 1(b) shows that embeddings of longer texts are more densely clustered near the origin in the dimensionally reduced embedding space, indicating a collapse that reduces variance among embeddings. Details of the rewriting process can be found in Appendix E.1. Figure 1(c) further demonstrates that embeddings of longer texts exhibit higher cosine similarity to each other, leading to smaller differences between them. This collapse results in a distributional inconsistency between embeddings of different text lengths, adversely affecting the performance of various downstream tasks (Yan et al., 2021) such as text classification.
@@
To study how text length affects the distribution of embeddings, we conduct a rigorous analysis of the self-attention mechanism in Fourier space (Wang et al., 2022b) (Section 2.2). Based on the finding that cascading self-attention blocks are equivalent to repeatedly applying a low-pass filter, we further prove that the attenuation rate of the low-pass filter is proportional to the largest singular value $\sigma_{a}$ of the high-frequency components (HC) in self-attention matrix. Furthermore, assuming that the input keys and query tokens follow a Gaussian distribution, we can prove that the $\sigma_{a}$ value decreases as text length increases. As a result, longer input texts retain more of the Direct Component (DC) in the token signals, which means that the embeddings at the output layer will collapse to a very narrow space. This theoretical proof explains the observations in Figure 1(b).
@@
Building on this theoretical analysis, we propose a simple yet effective solution named Temperature Scaling (TempScale) to mitigate the length collapse phenomenon. TempScale manipulates the calculated attention map by dividing the attention scores by a parameter named temperature smaller than 1 before applying the $\mathrm{softmax}(\cdot)$ operator. This adjustment increases the variance of the attention score matrix, as shown in Eqn. 3, leading to a larger $\sigma_{s}$ for the self-attention matrix. As a result, the self-attention matrix has a lower filter attenuation rate, leading to more diverse embeddings. As shown in Figure 1(c), a smaller temperature enables embeddings to exhibit lower pairwise cosine similarity and results in a more even distribution, thereby alleviating length collapse.
Our contributions can be summarized as follows:
  • We uncover the length collapse phenomenon and then establish a rigorous theoretical analysis from the spectral domain and show that the length collapse is due to the low-pass filtering strength of self-attention increases as the sequence length increases, leading the token signals retain only their DC component.
  • We present TempScale, a theoretically grounded scaling technique that incorporates a temperature parameter into the $\mathrm{softmax}(\cdot)$ function to achieve a higher attenuation rate in the self-attention matrix. TempScale is efficient, easy to use, tuning-free, and able to generalize across different embedding models compared to extending context window methods.
  • We conduct extensive experiments by integrating TempScale with mainstream embedding models. We demonstrate that TempScale can improve existing embedding models especially on long text inputs, bringing up to 0.53% performance gains on 40 different datasets from Massive Text Embedding Benchmark (MTEB) (Muennighoff et al., 2023) and 0.82% performance gains on 4 tasks from LongEmbed (Zhu et al., 2024).
2 Long Inputs Lead to Homogeneous Embeddings
@@
In this section, we first present our notations and define the problem. Then, we introduce the Transformer structure in the mainstream embedding model and briefly explain the Fourier transform used in (Wang et al., 2022b). Based on the Fourier transform, we show that the attention mechanism acts as a low-pass filter, and longer input sequences strengthen the filtering effect, leading to increasingly similar representations. This results in cosine similarity increasing with the length of the text.
@@ 2.1 Preliminaries and Background
Notations. Let $\bm{X}\in\mathbb{R}^{n\times d}$ denote the input feature matrix, where $n$ is the number of input tokens, and $d$ is the embedding dimension. Let $\bm{x}_{i}\in\mathbb{R}^{d}$ represent the vector corresponding to the $i$-th token and $\bm{z}_{j}\in\mathbb{R}^{n}$ represent the token sequence corresponding to the $j$-th dimension, where $i\in\{1,\dots,n\}$ denotes the $i$-th row, and $j\in\{1,\dots,d\}$ denotes the $j$-th column.
Transformer Architecture. In most modern embedding models (Chen et al., 2024; Xiong et al., 2021), a bidirectional transformer architecture based on attention mechanisms is widely used. These models generally consist of three key components: the embedding layer, a stack of transformer encoder blocks incorporating Multi-Head Self-Attention (MSA) and Feed-Forward Networks (FFN), and a pooling layer at the end to generate the final embedding representation of the input sequence. The Self-Attention(SA) module is the fundamental part of MSA, which takes inputs consisting of the token representations $\bm{X}$ from the previous layer, and it encodes each token by aggregating information from other tokens based on the attention scores, formulated as below (Vaswani, 2017):
$\operatorname{SA}(\bm{X})=\mathrm{softmax}\left(\frac{\bm{X}\bm{W}_{Q}(\bm{X}\bm{W}_{K})^{T}}{\sqrt{d}}\right)\bm{X}\bm{W}_{V},$ (1)
where $\bm{W}_{K}\in\mathbb{R}^{d\times d_{k}}$, $\bm{W}_{Q}\in\mathbb{R}^{d\times d_{q}}$, $\bm{W}_{V}\in\mathbb{R}^{d\times d}$ are the key, query and value weight matrices, respectively. The dimensions of the query and key vectors are denoted by $d_{q}$ and $d_{k}$, while $\sqrt{d}$ serves as a scaling factor to adjust the magnitude of the dot product. The function $\text{softmax}(\cdot)$ normalizes the attention scores row-wisely. Multi-Head Self-Attention (MSA) consists of SA heads, with their outputs combined through a linear projection:
@@
where the subscripts indicate the number of self-attention (SA) heads, $H$ represents the total number of heads, and $\bm{W}_{O}\in\mathbb{R}^{Hd\times d}$ projects the combined multi-head outputs back to the hidden dimension.
Fourier Analysis. We use Fourier transform as the main analytic tool in this paper as used in  (Wang et al., 2022b). Let $\mathcal{F}:\mathbb{R}^{n}\to\mathbb{C}^{n}$ represent the Discrete Fourier Transform (DFT), with its inverse, the Inverse Discrete Fourier Transform (IDFT), denoted by $\mathcal{F}^{-1}:\mathbb{C}^{n}\to\mathbb{R}^{n}$. Applying $\mathcal{F}$ to a token sequence $\bm{z}$ is equivalent to left multiplying a DFT matrix, where $k$-th row of DFT matrix denotes the Fourier basis corresponding to a certain frequency $\bm{f}_{k}=[\begin{array}[]{ll}e^{2\pi j(k-1)\cdot 0}&\ldots\quad e^{2\pi j(k-1)\cdot(n-1)}\end{array}]^{\top}/\sqrt{n}\in\mathbb{R}^{n}$, and $j$ is the imaginary unit. Let $\bm{\tilde{z}}=\mathcal{F}\bm{z}$ represent the spectrum of $\bm{z}$, where $\bm{\tilde{z}}_{dc}\in\mathbb{C}$ and $\bm{\tilde{z}}_{hc}\in\mathbb{C}^{n-1}$ correspond to the first element and the remaining elements of $\bm{\tilde{z}}$, respectively. We define the Direct-Current (DC) component of the input sequence $\bm{z}$ as $\mathcal{DC}[\bm{z}]=\bm{\tilde{z}}_{dc}\bm{f}_{1}\in\mathbb{C}^{n}$, and the complementary high-frequency component as $\mathcal{HC}[\bm{z}]=\left[\begin{array}[]{lll}\bm{f}_{2}&\cdots&\bm{f}_{n}\end{array}\right]\bm{\tilde{z}}_{hc}\in\mathbb{C}^{n}$, consistent with the definition in (Wang et al., 2022b).
In signal processing, a low-pass filter is a system that attenuates the high-frequency components of a signal while retaining the low-frequency components. In this paper, we specifically define a low-pass filter as one that preserves only the DC component $\mathcal{DC}\left[{\bm{z}}\right]$, while diminishing all other high-frequency components $\mathcal{HC}\left[{\bm{z}}\right]$. A more precise definition is provided in Definition 1.
@@
Let $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ be an endomorphism with $f^{t}$ obtained by applying $f$ for $t$ times. The function $f$ acts as a low-pass filter if and only if ${}\lim_{t\rightarrow\infty}\frac{\lVert\mathcal{HC}\left[{f^{t}(\mathbf{z})}\right]\rVert_{2}}{\lVert\mathcal{DC}\left[{f^{t}(\mathbf{z})}\right]\rVert_{2}}=0~{}$ for all $\bm{z}\in\mathbb{R}^{n}$.
For additional background information, please refer to Appendix A.
Figure 2: Visualization of the intensity of high-frequency components and their theoretical upper bounds. The blue line is defined by $\log(\lVert\mathcal{HC}\left[{\bm{X}_{l+1}}\right]\rVert_{F}/\lVert\mathcal{HC}\left[{\bm{X}_{l}}\right]\rVert_{F})$, and the red line is estimated using the results in Theorem 2. More details can be found in Appendix E.2.
@@ 2.2 Theoretical Analysis on Length Collapse
Overview. In this subsection, we aim to demonstrate that increasing the sequence length $n$ accelerates the rate of low-pass filtering in the attention matrix, leading to greater similarity in text embeddings for longer texts, and provide a theoretical justification of self-attention based on its spectral-domain effect. Based on the previous findings that self-attention is constantly a low-pass filter (Lemma 1), which continuously erases high-frequency information, our main result is that the largest singular value $\sigma_{a}$ of $\mathcal{HC}\left[{\bm{A}}\right]$ can influence the filtering rate, with a smaller $\sigma_{a}$ indicating that self-attention can eliminate more high-frequency information (Theorem 2). Furthermore, our analysis shows that longer input sequences lead to a smaller $\sigma_{a}$ value for the attention matrix, resulting in the embedding of longer texts losing more feature expressiveness (Theorem 3). In conclusion, based on the assumption that natural language texts exhibit relatively consistent means, we can infer that longer texts tend to yield more similar representations (Corollary 4) and cause the length collapse. Lastly, we discuss the impact of other modules in Transformer on length collapse and distinguish our work from prior studies on similar collapse phenomena induced by deep layers in Transformers.
Formally, the following lemma demonstrates that the attention matrix generated by a softmax function (e.g., Eqn. 1) acts as a low-pass filter, independent of the specific token features or context window.
@@
Lemma 1 follows directly from the Perron-Frobenius theorem (Meyer, 2000). Since all elements of the self-attention matrix are positive and each row sums to $1$, the largest eigenvalue is $1$. One can see that repeatedly applying the self-attention matrix can be viewed as the forward process of the embedding model. As the number of layers in the embedding model increases indefinitely, the final output retains only the DC component and thus loses all the feature expressive power. More detailed proofs of this lemma are provided in Theorem 1 in Wang et al. (2022b).
@@
(Filter Rate of SA) Let $\sigma_{a}$ be the largest singular value of $\mathcal{HC}\left[{\bm{A}}\right]$ and $\operatorname{SA}(\bm{X})=\bm{A}\bm{X}\bm{W}_{V}$ the output of a self-attention module, we have
@@
The proof of Theorem 2 can be found in Appendix B.1. Theorem 2 suggests the high-frequency intensity ratio to the pre- and post- attention aggregation is upper bounded by $\sigma_{a}\lVert\bm{W}_{V}\rVert_{2}$. When $\sigma_{a}\lVert\bm{W}_{V}\rVert_{2}<1$, $\mathcal{HC}\left[{\bm{X}}\right]$ converges to zero exponentially. We further present Figure 2 to justify our results, showing that the upper bound is consistent with the trend observed in the actual values.
Based on the proof that a lower $\sigma_{s}$ will lead to a higher filter-pass rate, we give proof that the $\sigma_{s}$ will decrease with input length $n$ increases in the following theorem.
@@
(Filter Rate of Different Input Length $n$) Let $\bm{XW}_{Q}$ and $\bm{XW}_{K}$ be a Gaussian matrix, where elements $q_{ij}\sim\mathcal{N}(0,\sigma_{q}^{2})$ and $k_{ij}\sim\mathcal{N}(0,\sigma_{k}^{2}),\forall i,j.$ Let $p_{ij}=\bm{q}_{i}^{\top}\bm{k}_{j}/\sqrt{d}$ the attention score of pair $i,j$, whose variance can be expressed as $\sigma_{s}^{2}=\sigma_{q}^{2}\sigma_{k}^{2}+C_{cross}$, where $C_{cross}$ is the cross-covariance of the squared queries and keys (Goodman, 1960). Then we have
@@
The proof of Theorem 3 is provided in Appendix B.2. Theorem 3 builds on the work of Fenton (1960), which addresses the sum of log-normal variables. As the input length $n$ increases, $\sigma_{a}$ decreases, leading to the suppression of more high-frequency information and a reduction in feature expressiveness due to the low-pass filtering effect of the self-attention matrix. To validate our conclusions, we sample texts of varying lengths and plot the $\sigma_{a}$ values of the attention matrix from the final layer of the model after inputting, as presented in Figure 5. The results indicate that $\sigma_{a}$ decreases as the text length increases, ultimately leading to a higher filtering rate.
To facilitate further analysis, we define the temperature of the SA as:
@@
(Length Collapse in Text Embeddings) Given two texts of length $n$, the cosine similarity of their text embeddings tends to increase as $n$ grows.
Figure 3: Mean pairwise cosine similarity of text embeddings across length intervals, with embeddings computed as the mean of token embeddings from the model’s word embedding matrix.
The proof of Corollary 4 can be found in Appendix B.3 and is primarily based on the assumption that the mean of word embeddings in natural language texts maintains a relatively consistent representation. Specifically, as illustrated in Figure 3, we get the text embedding by averaging the word embeddings derived from the BGE model’s embedding matrix for a given text. We then assess the similarity within specified length intervals. The results indicate that as text length increases, the similarity between any two embeddings also rises significantly, supporting the validity of the assumption. According to Theorem 3, with an increase in text length, text embeddings tend to converge more rapidly towards the consistent representation stated in our assumption, ultimately resulting in higher similarity for longer texts. Additionally, we analyze text embeddings for certain special sequences, such as repeated tokens. Our findings reveal that even when two texts have no overlapping tokens, the cosine similarity of their text embeddings approaches 1 as the sequence length increases. A comprehensive analysis of this phenomenon can be found in Appendix E.4. This further corroborates the existence of length collapse, even when the two sequences exhibit different mean word embeddings on average.

### v2

# Length-Induced Embedding Collapse in PLM-based Models

@@ Abstract
Text embeddings from PLM-based models enable a wide range of applications, yet their performance often degrades on longer texts. In this paper, we introduce a phenomenon we call Length Collapse, where embeddings of longer texts tend to cluster together. This clustering results in a distributional inconsistency between the embeddings of short and long texts. We further investigate how these differences contribute to the performance decline observed with longer texts across various downstream tasks. Through a rigorous theoretical analysis of the self-attention mechanism, which acts as a low-pass filter in PLM-based models, we demonstrate that as text length increases, the strength of low-pass filtering intensifies, causing embeddings to retain more low-frequency components. As a result, input token features become more similar, leading to clustering and ultimately the collapse of embeddings for longer texts. To address this issue, we propose a simple method, TempScale, which mitigates the Length Collapse phenomenon. By narrowing the gap in low-pass filtering rates between long and short texts, TempScale ensures more consistent embeddings across different text lengths. This approach leads to performance improvements of 0.94% on MTEB and 1.10% on LongEmbed, which focuses specifically on long-context retrieval, providing strong evidence for the validity of our analysis. The source code is available at https://github.com/Yuqi-Zhou/Length_Collapse.
Length-Induced Embedding Collapse in PLM-based Models
Yuqi Zhou, Sunhao Dai, Zhanshuo Cao, Xiao Zhang, Jun Xuthanks:   Corresponding author. Gaoling School of Artificial Intelligence, Renmin University of China, Beijing, China {yuqizhou, sunhaodai, caozhanshuo117, zhangx89, junxu}@ruc.edu.cn
@@ 1 Introduction
Text embeddings—dense vectors that capture the semantic information of texts—are essential for many NLP applications, including text analysis Aggarwal and Zhai (2012); Angelov (2020), question answering Tan et al. (2023); Xu et al. (2024), web search Zhao et al. (2024); Yates et al. (2021), and retrieval-augmented generation Gao et al. (2023); Fan et al. (2024). Typically, embeddings are generated by pre-trained language models (PLMs), which produce fixed-dimensional vectors regardless of text length. In practice, we expect PLMs to perform consistently across texts of varying lengths in downstream tasks.
Unfortunately, popular PLM-based embedding models perform poorly on longer texts. As shown in Figure 1(a), we evaluate mainstream models on the IMDB classification task from the Massive Text Embedding Benchmark (MTEB)Muennighoff et al. (2023), using test sets grouped by text length. The results show that models with varying capabilities degrade as text length increases. For example, the BGE Xiao et al. (2023) model’s accuracy drops from 75.6% in the [0, 100) token range to 59.0% in the [400, 500) range, a 16.6% point decline.
@@
We attribute this performance degradation to a systematic tendency in embedding models: as the text length $n$ increases, the cosine similarity between embeddings of different texts also increases—a phenomenon we refer to as Length Collapse. To verify this, we conduct controlled experiments shown in Figures 1(b) and 1(c). Figure 1(b) shows that embeddings of longer texts cluster more densely near the origin in the reduced embedding space, indicating a collapse that reduces variance. Figure 1(c) further shows that embeddings of longer texts have higher cosine similarity to each other, leading to smaller differences. This collapse causes distributional inconsistency between embeddings of different lengths. Furthermore, we analyze how this distributional inconsistency in embeddings of different text lengths leads to performance degradation in downstream tasks such as classification, retrieval, and STS, particularly affecting the performance of longer texts.
@@
To study how text length affects embedding distributions, we analyze the self-attention mechanism in Fourier space Wang et al. (2022b) (§ 2.2). Building on the findings that cascading self-attention blocks act as repeated low-pass filters, we prove that the attenuation rate of this filter is proportional to the largest singular value $\sigma_{a}$ of the high-frequency components in the self-attention matrix. Assuming Gaussian-distributed input keys and query tokens, we show that $\sigma_{a}$ decreases with increasing text length, causing longer texts to retain more of the Direct Component (DC) in token signals. This results in embeddings collapsing to a narrow space, as seen in Figure 1(b).
@@
To validate our theoretical analysis, we propose a method, Temperature Scaling (TempScale), to mitigate Length Collapse. TempScale manipulates the attention map by dividing the scores by a parameter called temperature (smaller than 1) before applying the $\mathrm{softmax}(\cdot)$ operator. This reduces the gap in $\sigma_{a}$ between long and short texts in the self-attention matrix, minimizing the distributional gap between their embeddings. As shown in Figure 1(c), a smaller temperature enables embeddings to exhibit lower pairwise cosine similarity, resulting in a more even distribution and alleviating Length Collapse. Additionally, TempScale improves performance by 0.94% and 1.10% in MTEB and LongEmbed Zhu et al. (2024), respectively. All results confirm the validity of our analysis.
Our contributions are as follows: ($1$) We are the first to identify Length Collapse, where embeddings of longer texts cluster together, and explain how this uneven distribution degrades overall performance, especially for long texts. ($2$) We provide a rigorous theoretical analysis in the spectral domain, demonstrating that Length Collapse occurs due to the increasing low-pass filtering strength of self-attention as sequence length grows, causing token signals to retain only their DC component. ($3$) To validate this theoretical analysis, we propose a simple method, TempScale, to mitigate Length Collapse. Empirical results show a 0.94% performance improvement across the MTEB and a 1.10% improvement on LongEmbed, validating the correctness of our theoretical analysis.
2 Length Collapse
@@
In this section, we define the problem and introduce our notations. We then present the Transformer structure in PLM-based models and briefly explain the Fourier transform from Wang et al. (2022b). Using this framework, we show that the attention mechanism functions as a low-pass filter, with longer input sequences amplifying this effect, resulting in increasingly similar representations.
@@ 2.1 Preliminaries and Background
Notations. Let $\bm{X}\in\mathbb{R}^{n\times d}$ denote the input feature matrix, where $n$ is the number of tokens and $d$ is the embedding dimension. Let $\bm{x}_{i}\in\mathbb{R}^{d}$ represent the vector for the $i$-th token, and $\bm{z}_{j}\in\mathbb{R}^{n}$ represent the token sequence for the $j$-th dimension, where $i\in\{1,\dots,n\}$ and $j\in\{1,\dots,d\}$.
Transformer Architecture. Most modern embedding models Chen et al. (2024a); Xiong et al. (2021) use a bidirectional transformer architecture with attention mechanisms. These models typically consist of three components: an embedding layer, a stack of transformer encoder blocks with Multi-Head Self-Attention (MSA) and Feed-Forward Networks (FFN), and a pooling layer to generate the final embedding. The Self-Attention (SA) module encodes each token by aggregating information from other tokens based on attention scores, as defined by the equation Waswani et al. (2017):
$\operatorname{SA}(\bm{X})=\mathrm{softmax}\left(\frac{\bm{X}\bm{W}_{Q}(\bm{X}\bm{W}_{K})^{T}}{\sqrt{d}}\right)\bm{X}\bm{W}_{V},$
where $\bm{W}_{K}$, $\bm{W}_{Q}$, and $\bm{W}_{V}$ are the key, query, and value weight matrices. The function $\text{softmax}(\cdot)$ normalizes the attention scores. Multi-Head Self-Attention (MSA) aggregates the outputs of multiple SA heads, projected back to the hidden dimension:
@@
where $H$ is the number of heads, and $\bm{W}_{O}$ projects the combined outputs to the hidden dimension.
Fourier Analysis. Inspired by the recent work Wang et al. (2022b), which reveals that self-attention in Vision Transformers acts as a low-pass filter and causes feature degradation when scaling up depth, we employ Fourier analysis to further explore this phenomenon. We use the Fourier transform as the primary analytic tool, following Wang et al. (2022b). Let $\mathcal{F}:\mathbb{R}^{n}\to\mathbb{C}^{n}$ represent the Discrete Fourier Transform (DFT), with its inverse $\mathcal{F}^{-1}:\mathbb{C}^{n}\to\mathbb{R}^{n}$. Applying $\mathcal{F}$ to a token sequence $\bm{z}$ corresponds to multiplying by a DFT matrix, where the $k$-th row corresponds to the Fourier basis at frequency $\bm{f}k=[e^{2\pi j(k-1)\cdot 0},\ldots,e^{2\pi j(k-1)\cdot(n-1)}]^{\top}/\sqrt{n}\in\mathbb{R}^{n}$, with $j$ as the imaginary unit. Let $\bm{\tilde{z}}=\mathcal{F}\bm{z}$ denote the spectrum of $\bm{z}$, with $\bm{\tilde{z}}_{dc}\in\mathbb{C}$ and $\bm{\tilde{z}}_{hc}\in\mathbb{C}^{n-1}$ as the DC and high-frequency components, respectively. We define the Direct-Current (DC) component as $\mathcal{DC}[\bm{z}]=\bm{\tilde{z}}_{dc}\bm{f}_{1}\in\mathbb{C}^{n}$, and the high-frequency component as $\mathcal{HC}[\bm{z}]=[\bm{f}_{2},\dots,\bm{f}n]\bm{\tilde{z}}{hc}\in\mathbb{C}^{n}$. In signal processing, a low-pass filter preserves low-frequency components while attenuating high frequencies. In this paper, we define a low-pass filter that retains only the DC component $\mathcal{DC}\left[{\bm{z}}\right]$, while diminishing the high-frequency components $\mathcal{HC}\left[{\bm{z}}\right]$. A precise definition is given in Definition 1.
@@
Let $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ be an endomorphism with $f^{t}$ denoting $f$ applied $t$ times. The function $f$ acts as a low-pass filter if and only if $\lim_{t\rightarrow\infty}\frac{\lVert\mathcal{HC}\left[{f^{t}(\mathbf{z})}\right]\rVert_{2}}{\lVert\mathcal{DC}\left[{f^{t}(\mathbf{z})}\right]\rVert_{2}}=0$ for all $\bm{z}\in\mathbb{R}^{n}$.
For additional details, refer to Appendix A.
@@ 2.2 Theoretical Analysis on Length Collapse
Overview. This subsection demonstrates that increasing text length $n$ accelerates low-pass filtering in the attention matrix, making text embeddings for longer texts more similar. We justify this by analyzing self-attention’s spectral-domain effect. Building on Lemma1, we show that the largest singular value $\sigma_{a}$ of $\mathcal{HC}\left[{\bm{A}}\right]$ influences the filtering rate, with a smaller $\sigma_{a}$ indicating greater high-frequency loss (Theorem2). Our analysis further reveals that longer texts result in smaller $\sigma_{a}$ values, causing longer text embeddings to lose feature expressiveness (Theorem3). Consequently, we infer that longer texts yield more similar representations (Corollary4), leading to Length Collapse.
Formally, the following lemma demonstrates that the attention matrix generated by a softmax function acts as a low-pass filter, independent of the specific token features or context window.
@@
Lemma 1 follows from the Perron-Frobenius theorem Meyer (2000). Since all elements of the self-attention matrix are positive and each row sums to $1$, the largest eigenvalue is $1$. Repeated application of the self-attention matrix represents the forward process of the embedding model, and as the number of layers increases, the output retains only the DC component. A more detailed proof can be found in Theorem $1$ of Wang et al. (2022b).
@@
(Filter Rate of SA) Let $\sigma_{a}$ be the largest singular value of $\mathcal{HC}\left[{\bm{A}}\right]$, and $\operatorname{SA}(\bm{X})=\bm{A}\bm{X}\bm{W}_{V}$ the self-attention output. We have:
@@
Theorem 2 suggests the high-frequency intensity ratio to the pre- and post- attention aggregation is upper bounded by $\sigma_{a}\lVert\bm{W}_{V}\rVert_{2}$. When $\sigma_{a}\lVert\bm{W}_{V}\rVert_{2}<1$, $\mathcal{HC}\left[{\bm{X}}\right]$ converges to zero exponentially. We further present Figure 6 in Appendix C.2 to justify our results, showing that the upper bound is consistent with the trend observed in the actual values. The proof of Theorem 2 can be found in Appendix B.1.
Based on the fact that a lower $\sigma_{s}$ leads to a higher filter-pass rate, we prove in the following theorem that $\sigma_{s}$ decreases as the input length $n$ increases.
@@
(Filter Rate of Different Input Length $n$) Let $\bm{XW}_{Q}$ and $\bm{XW}_{K}$ be a Gaussian matrix, where elements $q_{ij}\sim\mathcal{N}(0,\sigma_{q}^{2})$ and $k_{ij}\sim\mathcal{N}(0,\sigma_{k}^{2}),\forall i,j.$ Let $p_{ij}=\bm{q}_{i}^{\top}\bm{k}_{j}/\sqrt{d}$ the attention score of pair $i,j$, whose variance can be expressed as $\sigma_{s}^{2}=\sigma_{q}^{2}\sigma_{k}^{2}+C_{cross}$, where $C_{cross}$ is the cross-covariance of the squared queries and keys Goodman (1960). Then we have
@@
The proof of Theorem 3 is in Appendix B.2. Building on the work of Fenton (1960); Nahshan et al. (2024) on log-normal variables, Theorem 3 shows that as input length $n$ increases, $\sigma_{a}$ decreases, suppressing high-frequency information and reducing feature expressiveness due to the self-attention matrix’s low-pass filtering effect. To validate this, we sample texts of varying lengths and plot the $\sigma_{a}$ values from the final layer’s attention matrix, as shown in Figure 7 in Appendix C.3. The results confirm that $\sigma_{a}$ decreases with text length, leading to a higher filtering rate. To facilitate further analysis, we define the temperature of the SA defined in Nahshan et al. (2024) as:
@@
Figure 2: Mean cosine similarity of text embeddings across length intervals, with embeddings averaged from the model’s word embedding matrix.
@@
(Length Collapse in Text Embeddings) As the text length $n$ increases, the cosine similarity of their embeddings tends to increase.
The proof of Corollary 4 in Appendix B.3 is based on the assumption that the mean of word embeddings in natural language texts remains consistent, with the formal definition and proof detailed therein. As shown in Figure 2, we compute text embeddings by averaging word embeddings from the BGE model’s embedding matrix and then assess similarity across length intervals. The results show that as text length increases, embedding similarity rises, supporting the assumption. However, we must emphasize that the increased low-pass filtering with length is the primary cause of Length Collapse. In Appendix C.4, we demonstrate that repeating two different tokens and computing the similarity between the sequences shows that the cosine similarity of their embeddings increases with length, even without any overlapping tokens.
(a) Classification and Clustering
(b) Retrieval
(c) STS
Figure 3: A 3D toy example explains why Length Collapse leads to performance degradation.